OpenStudy (anonymous):

Currently, Sam's coach is twice as old as Sam. When Sam started soccer 10 years ago, his coach was four times his age. How old is Sam now?

OpenStudy (anonymous):

Are you doing FLVS algebra 1? O.o

OpenStudy (anonymous):

Let sam's coach's age be 2x so sam's age is x 4(x - 10) = 2x - 10

OpenStudy (ash2326):

let sam's age be x and his coach's age be y y=2x ten years ago sam's age is x-10 his coach's age is y-10 y-10=4(x-10) y=2x putiing in above equation 2x-10=4x-40 -2x=-30 x=15 y=30 sam's age is 15

OpenStudy (anonymous):

4x - 40 = 2x -10 2x = 30 x = 15 2x = 30